OpenStudy (anonymous):

What is the surface area of the prism? A. 82 yd2 B. 90 yd2 C. 96 yd2 D. 108 yd2

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

So the trick to right angle prisms is to just do it as if it were a rectangular prism and then divide the solution by 2.

OpenStudy (anonymous):

???????????

OpenStudy (anonymous):

So treat it as if it were a rectangular prism. I'll draw out the triangle side so you can understand where I'm coming from.

OpenStudy (anonymous):

When you solve for the area of a triangle, you're actually solving for the area of a hypothetical rectangle and then halving it to get the triangle. We're using the same concept for the surface area.

OpenStudy (anonymous):

Here is the triangle side of the figure you have. I drew it out and then doubled it to make a rectangle. I'd like you to solve for the area and perimeter of the drawing I attached.

OpenStudy (anonymous):

72?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Ok, so let's go through this step by step. First thing to do is solve this: Rectangular prism with length of 4, width of 3, and height of 5. Can you solve this?

OpenStudy (anonymous):

What si the formula again lol is it 2*b+P*H

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

Yes, that is the formula

OpenStudy (anonymous):

2*12+20*5

OpenStudy (anonymous):

The perimeter would just be 2(4) + 2(3) = 14 so 2*12 + 14*7. (The height should be 7 not 5, mess up on my part there)

OpenStudy (anonymous):

Ok so the answer is 122/2=61

OpenStudy (anonymous):

Ok, almost there. 61 is the answer of the surface area except for the hypotenuse are. So just find the length of the hypotenuse and multiply it by the length of the rectangle.

OpenStudy (anonymous):

What the!?! what is the hypotenuse???

OpenStudy (anonymous):

I dont get it! UGH I CANT DO THIS!

OpenStudy (anonymous):

Look at the figure. The hypotenuse should be 5.

OpenStudy (anonymous):

5*5=10 so 10

OpenStudy (anonymous):

5*length of rectangle. The length of the rectangle is 7 so its 5*7=35. Now add 61 to 35 to get the answer.

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

yes, the answer is C